OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. Round to tenths if necessary. 3x2 - 4x - 8 = 0 A. {-1.7, 2.8} B. {-2.8, 2} C. {-1.1, 2.4} D. no real solution

OpenStudy (hunus):

To find the roots of an equation of the form \[\Large ax^2 + bx + c = 0\]You can either factor it, or use the equation \[\Large x= \frac{-b \pm \sqrt{(b^2-4ac})}{2a}\]

OpenStudy (hunus):

So for your function, 3x^2 - 4x -8 = 0 what are the values of a, b, and c?

OpenStudy (anonymous):

whats the answer?

OpenStudy (hunus):

I'm not going to just tell you the answer

OpenStudy (hunus):

But I will help you solve it